### v1

Abstract
LLM agents excel in settings where the environment is largely static and the relevant information fits within the model’s context window, but struggle in open-ended investigations where explanations must be derived by iteratively mining evidence from massive, heterogeneous raw data. Such settings often contain hidden dependency structure: entities interact, signals co-vary, and the relevance of a detail may only become apparent after additional discoveries elsewhere. Bounded context window further forces agents to summarize and compress intermediate findings before their significance is known, increasing the chance that crucial evidence is discarded prematurely.
ReAct-style LLM agents are particulary brittle here for two structural reasons. First, they operationalize investigations as a linear retrieve $\rightarrow$  summarize $\rightarrow$ reason loop in which each step is effectively a conditional query on the current (partial) evidence set; this makes outcomes highly sensitive to exploration order in iterative investigation and discovery tasks, introduces run-to-run non-determinism, and creates a reliability gap where Pass@k can be high while Majority@k remains much lower, indicating success by chance rather than consistent reasoning. In these settings, simply scaling inference (e.g., sampling more rollouts or longer chains) does not reliably stabilize conclusions, since candidate explanations cannot be checked or grounded autonomously. As new evidence arrives, earlier interpretations must be revised, but ReAct provides no explicit mechanism for belief bookkeeping and revision. Second, ReAct entangles semantic reasoning with controller responsibilities (tool orchestration, state tracking, and execution hygiene), so execution failures (e.g., argument errors, repetition, plan drift) leak into and degrade the reasoning process while wasting scarce context budget.
[…]
footnotetext: 1IBM Research, Yorktown Heights, New York, USA. Correspondence to: Saurabh Jha <Saurabh.Jha@ibm.com>.
1 Introduction
Large language model (LLM) agents have demonstrated strong performance on structured reasoning benchmarks—code synthesis, mathematical proof, and symbolic planning—yet remain brittle on open-ended queries where explanations must be inferred from massive heterogeneous data. This challenge arises ubiquitously across domains, including medical diagnosis, forensic investigations, and modern IT operations. In the IT domain in particular—spanning SRE, FinOps, and DevOps—cloud infrastructure, microservice architectures, CI/CD pipelines, financial platforms, and security monitoring systems continuously generate large volumes of observability data, including logs, metrics, distributed traces, configuration states, and events. When anomalies occur—service degradations, budget overruns, security incidents, or release failures—operators must navigate this ocean of data to answer a deceptively simple question: what explains the observed behavior of the system? Answering such questions is at the heart of operational diagnosis.
Formally, operational diagnosis can be cast as an abductive reasoning problem over graphs. Given a set of observed symptoms (alerts, events, failures) from a dynamic environment—which are often partial, noisy, and evolving—the goal is to infer a concise and coherent explanation that accounts for these observations in terms of the underlying system structure and dynamics.
Limits of existing agent patterns:
Current agentic approaches struggle to automate operational diagnosis. While the limitations of ReAct-style agents—such as tool invocation failures and redundant or degenerate tool calls—are well known Cemri et al. (2025), the dominant challenge in operational diagnostics lies in context gathering and sequencing under uncertainty. Reasoning is only as sound as the evidence it conditions on: if relevant context is not retrieved, it is ignored and cannot be reasoned over, regardless of model capacity. In open-ended investigations, however, the relevance of evidence is not known a priori, and ReAct agents lack principled mechanisms for ensuring systematic context acquisition. Tool calls are issued opportunistically, leading to missed low-frequency but high-impact signals or excessive retrieval of irrelevant data. A common mitigation is to summarize retrieved evidence before further reasoning, but this introduces a fundamental failure mode: summarization necessarily discards information before its relevance can be established, causing critical evidence to be lost prematurely Liu et al. (2024).
More fundamentally, operational diagnosis exhibits non-monotonic belief revision behaviors: the interpretation of an entity or signal naturally changes as new evidence is discovered elsewhere in the system, making reasoning order-dependent. As a result, the sequence in which evidence is retrieved, summarized, and presented to the LLM materially affects conclusions. In practice, ReAct agents exhibit high variance across runs due to planning non-determinism, unstable exploration orderings, and lossy intermediate summarization, producing explanations that are incomplete, irreproducible, and fragile. These failures are architectural: ReAct implicitly assumes linear, monotonic reasoning over an accumulating context, whereas operational diagnostics require structured context acquisition with explicit support for ordering and belief revision.
Towards graph-guided agentic explanations:
Building on the abductive framing above, we introduce an operational-graph abstraction for open-ended investigations under partial observability. Across IT domains such as reliability, cost management, security, and deployment, abductive explanations are naturally structured by dependency relationships among system entities. This motivates representing the environment as an operational graph and defining diagnosis as explanations over graphs: the agent reasons over a partial, evolving world model, while explanations correspond to minimal subgraphs consistent with observed evidence. Framing diagnosis in this way converts an ill-defined exploration problem into structured inference over a world model.
Crucially, diagnosis is inherently non-monotonic—beliefs about the explanatory role of an entity may change as new evidence is uncovered elsewhere in the graph. Drawing inspiration from belief propagation Pearl (1988), we model diagnosis as an iterative process in which local beliefs are assigned, revised, and propagated through the graph. This perspective naturally leads to a disaggregated (neurosymbolic) architecture in which the language model performs local reasoning over bounded neighborhoods, while a symbolic controller manages traversal order, context gathering, belief bookkeeping, and termination. This separation is essential for supporting non-monotonic belief revision: beliefs are recorded explicitly and revised as new evidence is incorporated, rather than being implicitly overwritten in a growing prompt. By externalizing planning and context management from the LLM, the system keeps context bounded, greatly reduces tool-use failures, and produces explanations that are reproducible and auditable. This separation also addresses production constraints: bounded per-node context scales to large observability datasets, while deterministic control logic minimizes LLM invocations, reducing exposure to API rate limits.
Contributions:
While addressing limitations of existing agent patterns, this paper makes the following contributions:
  • Separation of concerns between deterministic and non-deterministic reasoning: By casting operational diagnosis as structured inference over an explicit operational graph, we shift diagnosis from free-form conversational reasoning tackled by non-deterministic LLMs to the discovery of minimal explanatory subgraphs consistent with observed evidence. This leads to a novel agentic architecture where deterministic and non-deterministic reasoning work in tandem. More specifically, we derive a disaggregated agent architecture in which language models perform bounded local reasoning, while a symbolic controller manages traversal order, context gathering, belief bookkeeping, and termination. This separation enables bounded context, deterministic control flow, reduced tool-use failures, and explicit belief revision.
  • Belief-revision perspective on agentic diagnostics. We identify non-monotonic belief revision as the central challenge in LLM-based diagnostics and show that explanation quality, stability, and ordering are fundamentally coupled through this lens. We connect failures of existing agent patterns to their inability to explicitly represent and revise beliefs under evolving evidence. Drawing inspiration from belief propagation, we introduce a principled agent model in which diagnosis proceeds via iterative local belief assignment and propagation over an operational graph. This model provides a unifying abstraction for context acquisition, reasoning, and explanation in dynamic environments.
  • Validation on operational diagnosis tasks. We validate the proposed formulation and agent model on SRE diagnosis tasks using ITBench Jha et al. (2025), demonstrating improvements in accuracy, stability, and token efficiency over ReAct-style agents. While this paper focuses on operational diagnostics, the same principles have been applied to software bug localization in production systems, in our concurrent work Cui et al. (2025).
2 Problem Formulation
2.1 The Explanations-over-Graphs (EoG) Problem
We model the environment as an Operational Graph $G=(V,E)$ representing a system topology. In the IT domain, this graph may represent various IT specific components such as applications, services, resources, and code modules. In general, Nodes $V$ represent entities; edges $E$ capture structural relationships (calls, depends-on, contains). Crucially, $E$ encodes topology, not the direction of fault propagation—that must be inferred.
Alerts are events that maps directly to nodes $G$ and we do track them accordingly. Given a set of alerts $A\subseteq V$ (entities exhibiting anomalous behavior), the goal is to discover:
  1. 1.
    An Explanatory Subgraph $G_{S}=(V_{S},E_{S})$ where $V_{S}\subseteq V$ is the relevant active set and $E_{S}$ are directed edges representing inferred propagation explaining the fault. We use the notation $u\to v$ to mean that “$u$ explains vv”.
  2. 2.
    A Frontier $\mathcal{F}\subseteq V_{S}$: the minimal set of origin entities that explain all alerts in $A$.
We use “explanatory” rather than “causal” deliberately: the goal is to infer $G_{S}$ from observational evidence (logs, traces) rather than through interventions. The resulting graph represents the system’s explanation for the observed behavior.
3 Characterizing failure modes of ReAct Agent critical for Operational Diagnostic Tasks
We motivate our approach by characterizing the failure modes of ReAct agents Yao et al. (2023) on ITBench Jha et al. (2025), a benchmark of operational diagnostic tasks. Each ITBench scenario presents an incident in a Kubernetes environment—a service degradation, misconfiguration, or resource failure—and requires the agent to identify the originating cause by querying logs, metrics, traces, and configurations. Scenarios are annotated with ground-truth (GT) entities (the originating cause), non-GT entities (components exhibiting symptoms), and the evidence (log patterns, metric anomalies, error messages) that connects symptoms to their cause. This structure enables post-hoc analysis of where agents fail in the diagnostic pipeline. While we focus on ITBench in this section, the same diagnostic structure—entities, evidence, and explanatory chains—applies to other operational domains, including identifying code defects in production systems Cui et al. (2025) and anomaly explanation in cloud cost management.
We evaluate a range of LLM models (including o4-mini, Gemini 2.5 Pro/Flash, and Claude Opus 4.5) under the ReAct paradigm across multiple runs per task. Our analysis reveals that diagnostic failures stem not merely from lack of domain knowledge, but from structural deficiencies in the ReAct paradigm itself: stochasticity in exploration, brittleness in tool execution, and sensitivity to context ordering.
3.1 The Reliability Gap: Stochasticity vs. Reasoning
Figure 1: The reliability gap between Pass@$k$ (success in at least one run) and Majority@$k$ (success in the majority of runs). Large gaps indicate stochastic success rather than consistent reasoning.
We first examine absolute diagnostic performance. Across models, mean recall ranges widely: Gemini 3 Flash achieves the highest at 76%, followed by Gemini 3 Pro (50%) and Claude Opus 4.5 (53%). The remaining models fall below 35%, with GPT-OSS-120B at just 12%. These results indicate that, under the ReAct paradigm, most frontier models remain far from deployment-ready on operational diagnostic tasks.
O1:  Mean recall ranges from 12% to 76%. Most models are not deployment-ready.
However, even models with high task completion rates exhibit significant variance across runs. Figure 1 illustrates the gap between Pass@$k$ (success in at least one of $k$ runs) and Majority@$k$ (success in the majority of runs). Gemini 3 Flash achieves 94% Pass@$k$ but drops to 74% Majority@kk—a 20-point gap. Claude Opus 4.5 shows an even larger disparity: 74% Pass@$k$ versus 49% Majority@$k$, with only 34% of scenarios yielding consistent correct diagnoses across all runs. This reliability gap—increasingly recognized as a key challenge in production agentic systems Kartik et al. (2025)—indicates that correct diagnoses often result from favorable exploration paths rather than robust reasoning.
O2:  Large gap between Pass@$k$ and Majority@kk—models succeed by chance, not consistent reasoning.
3.2 Exploration Pathology: The Case for Graph Traversal
(a) GT discovery and assertion failures.
(b) Evidence discovery and use failures.
(c) Exploration vs recall.
Figure 2: Exploration pathology in ReAct agents. (a) Failure to reach the causal entity. (b) Failure to fetch relevant evidence. (c) Correlation between exploration breadth and success. These patterns support modeling diagnosis as iterative graph traversal with bounded context fetching at each hop.
Operational diagnosis is a search problem over a massive state space. Our analysis reveals that ReAct agents fail this search not because they cannot reason, but because they lack the structural priors to navigate system topology.
Origin entity (GT) discovery failures.
The first hurdle is reachability. Figure 2(a) tracks the GT entity through five mutually exclusive stages: GT never saw—the entity never appeared in any tool output (exploration failure); GT saw, not queried—the entity appeared passively but was never actively investigated (attention failure); GT queried, not asserted—the agent investigated but never identified it as the cause (reasoning failure); GT asserted, not in output—the agent identified the GT but omitted it from the final answer (prioritization failure, often due to insufficient evidence or noisy confounding signals); and GT asserted, in output—successful diagnosis. A substantial failure mode is “GT never saw”: in any diagnostic query over a graph—and distributed systems naturally enforce a topology graph—the origin entity is linked to the symptom via dependency edges, yet unstructured data discovery through the ReAct pattern fails to follow these links. Therefore, the diagnostic process must be modeled as a graph traversal, where the agent follows dependency edges from symptoms toward upstream causes.
Evidence discovery failures.
Figure 2(b) shows high incidence of Evidence not in context (red): the necessary logs or metrics existed but were never fetched. This failure mode arises from three distinct causes: (1) the entity containing the evidence was never explored, so the evidence was never seen; (2) tool arguments were too restrictive to retrieve the relevant data (e.g., fetching 5 log lines instead of 50 to conserve context budget); and (3) what constitutes “relevant” data may not be known a priori—an agent cannot know which log lines matter until later evidence reveals their significance (the sequencing problem in §3.4). Evidence ignored (orange) indicates evidence appeared but was never investigated—a reasoning failure. Evidence used (green) indicates successful retrieval and utilization. Notably, the “evidence ignored” category is relatively small for recent frontier models, indicating that modern LLMs understand IT diagnostic semantics and can utilize evidence when present; infusing additional domain knowledge is orthogonal to our contribution. Since agents cannot scan the whole system at once, they must iteratively explore each node’s neighborhood, fetching and analyzing local evidence hop-by-hop.
Structured exploration failures.
Figure 2(c) shows that success rates correlate strongly with exploration breadth: agents exploring fewer than 4 entities achieve below 20% success; those exploring 8+ entities exceed 60%. However, exhaustive exploration is infeasible in production systems with thousands to millions of entities. ReAct agents, which treat every tool call as an independent action, lack mechanisms to systematically expand the frontier of investigated nodes while following evidence. Guided exploration along dependency edges—rather than unstructured data exploration—addresses this gap.
O3:  High rates of “GT never saw” and “Evidence not in context” failures due to lack of structured exploration.
I1:  Enforce structured exploration via iterative graph traversal, guided by evidence.
3.3 Controller Brittleness: The Cost of Interleaved Execution
(a) Plan abandonment rate.
(b) Tool repetition rate.
(c) Tool failure categories.
Figure 3: Controller brittleness in ReAct agents. (a) Plan abandonment rate: frequency with which agents discard stated investigation plans. (b) Tool repetition rate: percentage of tool calls that duplicate previous invocations. (c) Tool failure categories: distribution of syntactic vs. semantic errors.
The ReAct paradigm conflates reasoning (what to investigate) with execution (how to invoke tools). This overloading of the LLM leads to measurable controller instability.
Plan abandonment.
Figure 3(a) quantifies how often agents abandon their stated investigation plans. o4-mini abandons plans in over 50% of cases; even stronger models like Claude Opus 4.5 exhibit abandonment rates over 50%. Plan abandonment occurs for three reasons: (1) agents fail to revise plans as new evidence emerges, leaving stale steps that no longer apply Gandhi et al. (2025); (2) agents forget earlier plan steps as context grows (context rot) Liu et al. (2024); and (3) agents prematurely conclude they have found the cause and skip remaining steps, converging to a minimal explanation before sufficient evidence is gathered. This instability undermines auditability: the agent’s stated reasoning diverges from its actual behavior.
Tool repetition.
Agents exhibit degenerate loops where they repeatedly invoke the same tool calls. Figure 3(b) shows that GPT-OSS-120B has a repetition rate of about 12%, consuming context budget without advancing the investigation. Tool repetition can occur for two reasons: (1) re-executing a failed tool call without correcting the arguments or function signature; or (2) re-fetching data to analyze it differently in light of new evidence. The latter is potentially useful—revisiting data with updated hypotheses is a valid diagnostic strategy—but we find no evidence of productive re-analysis in our trajectories; the vast majority of repetitions are failed-execution loops. This redundancy compounds with trajectory length, leaving less capacity for novel exploration.
Syntactic failures.
Figure 3(c) categorizes tool invocation failures. A substantial fraction are strictly syntactic—malformed arguments, type errors, invalid parameter combinations—rather than semantic reasoning failures Chen et al. (2024b). These errors indicate that LLMs are suboptimal drivers for precise tool orchestration when simultaneously reasoning about complex domains. In contrast, conventional non deterministic programs tend to excel in tool invocation and execution, when they have access to the right contextual information.
O4:  High rates of plan abandonment, tool repetition, and syntactic failures.
I2:  Separation of concerns—decouple Controller (syntax, state) from Reasoner (semantics).
3.4 Path Dependence and Order Sensitivity
Figure 4: Impact of oracle reordering (i.e., presenting relevant entities/evidence earlier in the exploration sequence) on majority@k.
We identify that ReAct agents are highly sensitive to the order in which information is discovered. In an oracle experiment (Figure 4), we reorder the investigation sequence of the agent’s original trajectory to prioritize the path from symptoms to ground-truth entities, without adding any new entities or information, and then re-execute with identical model and tools.
This reordering yields 14–70% improvement in Majority@$k$ across all models, with GPT-OSS-120B showing 70% gains and o4-mini showing 61%. The effect is consistent across model scales and families, indicating a structural property of the paradigm rather than model-specific limitations.
This path dependence confirms that agents struggle with non-monotonic reasoning: they form early beliefs based on initial context and fail to revise them when downstream evidence contradicts upstream assumptions. In diagnostic domains, where the relevance of evidence often becomes apparent only after subsequent discoveries, this rigidity is particularly costly.
O5:  Oracle reordering yields 14–70% improvement in Majority@kk—exploration order materially affects success.
I3:  Belief revision—new evidence must propagate backward to update prior judgments.
3.5 Summary: From Observations to Design Principles
The preceding analysis identifies three structural deficiencies in the ReAct paradigm that limit its efficacy in operational diagnosis:
  1. 1.
    Unguided exploration drives the reliability gap (O2) and exploration failures (O3). The correlation between entity coverage and success motivates I1: use graph topology to guide traversal and ensure systematic coverage.
  2. 2.
    Interleaved reasoning and execution cause controller brittleness (O4). High rates of syntactic failure and plan abandonment motivate I2: delegate tool orchestration and state management to a symbolic controller while restricting the LLM to semantic reasoning over a bounded context.
  3. 3.
    Path-dependent reasoning hinders belief revision (O5). The sensitivity of success rates to investigation order motivates I3: maintain an external state of local assessments that can be revised as new evidence emerges, enabling non-monotonic updates.
These failure modes reinforce one another. Controller brittleness consumes context that could otherwise support broader exploration; incomplete exploration leads to missed evidence; and rigid reasoning states prevent recovery when initial assumptions prove incorrect. The following section formalizes an architecture that addresses these deficiencies through the Explanations-over-Graphs framework.
4 Methodology
[…]
Figure 5: Semantic Belief Propagation architecture, illustrated with the worked example (§4.5). Left: The topology graph $G$ is incomplete—the $S_{2}\to S_{3}$ dependency is not registered; additional services ($S_{5}$$S_{7}$) are uninvolved (healthy). An alert on $S_{2}$ seeds the ActiveSet. Center: The active inference loop: pop from ActiveSet, fetch context, classify via policy, update state. New discoveries or belief changes add nodes back to ActiveSet. Right: The output explanatory graph $G_{S}$ with the discovered edge (dashed) and final beliefs. Origin nodes use double circles; symptom nodes use hatched fill. The frontier $\mathcal{F}=\{S_{1}\}$ identifies the flash sale as the true origin.
We propose Semantic Belief Propagation (SBP), an architecture designed to address the structural failures of ReAct agentic loops identified in §3.5: unguided exploration (O3, I1), controller brittleness (O4, I2), and path-dependent reasoning (O5, I3). We implement this architecture as the EoG (Explanations over Graph) agent, which performs distributed causal inference over operational graphs, in contrast with existing monolithic or hierarchical approaches typically relying on the LLM to simultaneously manage exploration, reasoning, and global state. This coupling makes the resulting existing systems fragile: a single planning error or context overflow causes the entire diagnosis to fail.
SBP decomposes this loop into a distributed message-passing algorithm inspired by classical belief propagation Pearl (1988), enforcing a strict separation of concerns:
  • State (The Explanatory Graph): Instead of relying on an ephemeral context window, we maintain a persistent Explanatory Graph $G_{S}$. This structure acts as explicit external memory, storing the evolving belief state (e.g., $S_{1}$ is an Origin, $S_{2}$ is a Symptom) separate from the static system topology—addressing I3.
  • Decision (The Policy): We restrict the LLM to a stateless Abductive Policy $\pi_{\text{abd}}$. By removing the burden of global planning, the policy can focus entirely on local inference—mapping bounded evidence (logs, neighbor messages) to causal roles—addressing I2.
  • Control (The Algorithm): We replace the unstable “LLM-as-Planner” with a Deterministic Controller. This symbolic engine offloads housekeeping duties (queue management, message routing, termination), enforcing a rigorous exploration strategy that prevents the agent from “getting lost” or abandoning the plan, addressing I1.
This architecture enables non-monotonic reasoning: as new evidence propagates (e.g., a downstream service reports it is a symptom), upstream nodes can revise their initial beliefs, correcting early misdiagnoses without restarting the investigation.
Handling Inconclusive Results.
The agent may not always produce a definitive diagnosis. Two failure modes are possible: (1) evidence absence—the relevant observability data does not exist or is not retrievable via the CxC (e.g., logs were not retained, metrics not instrumented); and (2) evidence blindness—the policy $\pi_{\text{abd}}$ lacks the domain knowledge to recognize available evidence as diagnostically relevant. In either case, nodes are labeled Defer. If no Origin is identified at termination (i.e., $\mathcal{F}=\emptyset$), the Controller falls back to a best-effort ranking of candidate nodes based on accumulated evidence, explicitly signaling uncertainty to the operator.
4.1 Architecture Overview
The architecture comprises three components (Figure 5):
The Deterministic Controller.
A symbolic engine that maintains the global Ledger, manages the ActiveSet (nodes to investigate), and routes messages. It executes the algorithm logic (termination, damping) without performing semantic reasoning. This separation ensures safety and reproducibility: the agent cannot execute arbitrary code or tools; it can only request data via the Context Contract, preventing harmful side effects. It also provides operational efficiency: graph traversal, message routing, and bookkeeping require no LLM calls, bounding inference costs and API utilization to $O(|V_{S}|)$ policy invocations.
The Context Contract (CxC).
A strict interface that bounds and focuses the information provided to the policy, drawing on principles of active context management Li et al. (2025); Zhang et al. (2025a). When investigating node $v$, the CxC fetches a Context Packet containing:
  1. 1.
    Local observability data (logs, metrics) within the incident time window.
  2. 2.
    Topological context (immediate neighbors).
  3. 3.
    Incoming messages from neighbors (the Inbox).
Relevance Filtering: Domain experts can equip the CxC with specialized tools (e.g., “fetch top-k log anomalies”) to minimize evidence omission. This ensures the policy sees the most relevant signal without context overflow.
The Abductive Policy ($\pi_{\text{abd}}$).
A stateless, LLM-parameterized policy that performs local abductive inference. Given a bounded Context Packet for node $v$, $\pi_{\text{abd}}$ produces: (i) a local belief $B_{v}$, (ii) propagation claims $\mathcal{P}_{v}$, and (iii) a set of candidate entities $\mathcal{N}_{next}$ for further investigation. While the Controller governs execution and termination, $\pi_{\text{abd}}$ determines the epistemic trajectory of the search. We use “policy” in the decision-theoretic sense: a mapping from local observations to epistemic decisions (belief assignment and hypothesis selection), not as an online reward-optimizing controller.
4.2 State Formalism
We distinguish between three levels of state:
1. The Global Ledger ($\mathcal{L}$).
The immutable history of all reasoning steps, serving as an external scratchpad Nye et al. (2021) that enables replay and debugging.
2. Local Belief ($B_{v}$).
The current assessment of a node $v$, derived from the latest policy execution: $B_{v}=(L_{v},\mathcal{E}_{v})$ where:
  • $L_{v}\in\{\textsc{Healthy},\textsc{Origin},\textsc{Symptom},\textsc{Defer}\}$ is the explanatory label:
    • Healthy: Node is operating normally within the incident window.
    • Origin: Node has a citable change (config, deployment, resource) that explains the incident.
    • Symptom: Node is degraded, but evidence points to an upstream cause.
    • Defer: Evidence is inconclusive—either absent or unrecognized by the policy.
  • $\mathcal{E}_{v}$ is the evidence summary.
3. Messages ($m_{u\to v}$).
A mechanism to transport beliefs. A message contains a snapshot of $B_{u}$ at time $t$. When node $v$ processes its Inbox, it reads these messages to reason about upstream dependencies (the Local Markov Property).
4.3 Algorithm: Semantic Belief Propagation
We present a unified algorithm that interleaves exploration (finding new nodes) and propagation (refining beliefs). The system is implemented as an actor model, allowing nodes to process messages concurrently.
4.3.1 Initialization (Bootstrap)
The Controller initializes the investigation:
  • Define time window $W$ based on incident timestamp.
  • Initialize $\text{ActiveSet}\leftarrow A$ (or other domain heuristics).
  • Initialize $V_{S}\leftarrow\emptyset,E_{S}\leftarrow\emptyset$.
  • Set $\text{Budget}\leftarrow\text{max number of investigation hopes}$
4.3.2 Active Inference Loop
The core loop processes nodes in the ActiveSet until convergence (no new messages) or budget exhaustion (tracking the maximum number of investigation hops).
1:while $\text{ActiveSet}\neq\emptyset\land\text{Budget}>0$ do
2:  $v\leftarrow\text{ActiveSet}.\textsc{Pop}()$
3:  $V_{S}.\text{add}(v)$
4:  // 1. Fetch Context
5:  $\text{Inbox}\leftarrow\text{CollectMessages}(v)$
6:  $\text{Packet}\leftarrow\text{CxC.Fetch}(v,W,\text{Inbox})$
7:  // 2. Execute Policy
8:  $(B_{v}^{new},\mathcal{P}_{v},\mathcal{N}_{next})\leftarrow\pi_{\text{abd}}(\text{Packet})$
9:  // 3. Damping (Heuristic)
10:  if $\textsc{Flips}(v)>k_{\text{thresh}}$ then
11:   $B_{v}^{new}.L\leftarrow\textsc{Defer}$
12:  end if
13:  // 4. Update State
14:  $\text{Ledger}.\text{append}(v,B_{v}^{new})$
15:  for $(u\to v)\in\mathcal{P}_{v}$ do
16:   $E_{S}.\text{add}(u\to v)$
17:   if $u\notin V_{S}$ then
18:     $\text{ActiveSet}.\text{add}(u)$ $\triangleright$ Exploration
19:   end if
20:  end for
21:  // 5. Propagation
22:  if $B_{v}^{new}\neq B_{v}^{old}$ then
23:   $\text{Broadcast}(B_{v}^{new}\text{ to }\mathcal{N}(v))$
24:   $\text{ActiveSet}.\text{add}(\mathcal{N}(v))$ $\triangleright$ Re-activate neighbors
25:  end if
26:  // 6. Propose (Optional)
27:  for $u\in\mathcal{N}_{next}$ do
28:   if $u\notin V_{S}$ then
29:     $\text{ActiveSet}.\text{add}(u)$
30:   end if
31:  end for
32:  $\text{Budget}$-- $\triangleright$ Decrement Budget
33:end while
34:// Post-processing: Compute Frontier
[…]
The loop terminates when no more messages are pending (fixed point) or budget is exhausted. The Frontier $\mathcal{F}$ is computed after convergence as a post-processing step over the final belief state.
4.4 Properties and Correctness
Minimality and Correctness.
The algorithm seeks a Frontier $\mathcal{F}$ of irreducible origins.
$\mathcal{F}=\{v\in V_{S}:$ $L_{v}=\textsc{Origin}\land\nexists u\in V_{S}$ (1)
$\text{s.t. }L_{u}=\textsc{Origin}\land u\rightsquigarrow_{E_{S}}v\}$
By definition, $\mathcal{F}$ is minimal relative to the discovered graph $G_{S}$. If no node is labeled Origin (i.e., $\mathcal{F}=\emptyset$), the finalize phase falls back to best-effort ranking of candidates based on accumulated evidence. The correctness relies on three assumptions:
  1. 1.
    Completeness: Relevant data exists and is retrievable via CxC.
  2. 2.
    Soundness: The policy $\pi_{\text{abd}}$ correctly infers local dependencies.
  3. 3.
    Connectivity: Origins are reachable from alerts via the topology or discovered edges.
If these hold, the message passing ensures that evidence from origins propagates to symptoms, correcting any initial misclassifications caused by exploration order.
Termination.
Termination is guaranteed by the damping heuristic. If a node oscillates labels more than $k_{\text{thresh}}$ times, it is forced to Defer (an absorbing state). This bounds the total number of inference steps to $O(k_{\text{thresh}}\cdot|V_{S}|)$. Cycles in the topology are allowed, but the explanatory graph $E_{S}$ is directed by evidence. When evidence is conflicting, oscillation can occur; damping converts this to the absorbing Defer state and guarantees a fixed point when no messages remain.
4.5 Worked IT Example: Cascading Belief Revision
[…]
The diagnostic trap.
$S_{4}$’s Out Of Memory (OOM) error is strong local evidence of being the origin—memory exhaustion is a classic root cause. Without belief messages, the algorithm would terminate with $S_{4}$ as origin, leading operators to restart $S_{4}$ (which doesn’t address the upstream traffic pattern). This example shows both edge discovery (finding the hidden $S_{2}\to S_{3}$ dependency) and belief revision (correcting the $S_{4}$ misdiagnosis).
[…]
Figure 6: Edge discovery and belief revision. Row 1: The hidden edge $S_{3}\to S_{2}$ (dashed) is discovered via HTTP 500 traces—it was not in the topology graph. $S_{4}$’s OOM is the sole origin ($\mathcal{F}=\{S_{4}\}$), but traffic analysis reveals elevated load; $S_{4}$ revises and broadcasts. Row 2: Messages cascade upstream; $S_{1}$ (flash sale) is identified as the true origin. Origin nodes use double circles; symptom nodes use hatched fill.
  1. 1.
    $t=1$, Process $S_{2}$ (Alg. Steps 1–2, 4, 6): Alert node. CxC returns HTTP 500 errors showing failed calls to $S_{3}$. Edge discovery: $S_{2}\to S_{3}$ was not in the topology graph, but traces reveal this dependency. The policy adds edge $S_{3}\to S_{2}$ to $E_{S}$ and proposes $S_{3}$. $B_{S_{2}}=\textsc{Symptom}$ of $S_{3}$.
  2. 2.
    $t=2$, Process $S_{3}$: Logs show timeouts calling $S_{4}$. Proposes $S_{4}$. $B_{S_{3}}=\textsc{Symptom}$ of $S_{4}$.
  3. 3.
    $t=3$, Process $S_{4}$: Logs show OutOfMemoryError. Strong local evidence of being the origin—memory exhaustion is a classic root cause. $B_{S_{4}}=\textsc{Origin}$.
    At this point: $\mathcal{F}=\{S_{4}\}$. Single origin, plausible diagnosis. Without belief messages, we would stop here and blame $S_{4}$’s memory issue.
  4. 4.
    $t=4$, $S_{4}$ re-enters ActiveSet (Alg. Step 5: Propagation): CxC now includes traffic metrics (not just error logs). Policy sees: incoming request rate from $S_{3}$ is elevated (+30% above baseline). Re-reasons: “My OOM was caused by traffic overload, not internal failure.” $B_{S_{4}}\to\textsc{Symptom}$ of $S_{3}$. Broadcasts to $S_{3}$.
  5. 5.
    $t=5$, $S_{3}$ receives message (Alg. Step 5: Propagation continues): Inbox contains $B_{S_{4}}$: “I failed from your traffic.” $S_{3}$ checks its own incoming traffic—+25% from $S_{2}$. Propagates message upstream. Broadcasts to $S_{2}$.
  6. 6.
    $t=6$, $S_{2}$ receives message: Inbox contains $B_{S_{3}}$: “The cascade started from your traffic.” $S_{2}$ checks incoming—+20% from $S_{1}$. Proposes $S_{1}$.
  7. 7.
    $t=7$, Process $S_{1}$: Logs show flash sale triggered traffic spike. $B_{S_{1}}=\textsc{Origin}$.
  8. 8.
    Convergence (Alg. Post-processing): $\mathcal{F}=\{S_{1}\}$—correct. The flash sale is the true origin.
Key insight.
[…]
5 Implementation & Results
5.1 Implementation
We implement the EoG agent and a baseline ReAct agent on top of the Codex agent framework. [^1] Available at https://github.com/openai/codex The implementation comprises approximately 5,000 lines of Rust code. The Deterministic Controller manages the ActiveSet and Ledger with read-write locks, enabling concurrent node processing while ensuring atomic belief updates. The actor-model implementation supports asynchronous execution without corrupting global state even when messages arrive out of order. Observability data is retrieved through the Model Context Protocol (MCP), a standardized interface for LLM tool use. We implement domain-specific MCP servers exposing the operational graph and observability signals—including distributed traces Kaldor et al. (2017)—via the tools in Table 1.
Tool Description
alert_summary Active alerts with severity
get_context_contract Context Packet for entity
topology_analysis Dependency graph queries
metric_analysis Metrics with anomaly detection
log_analysis Service logs with filtering
event_analysis Platform events in time window
get_trace_error_tree Trace error propagation
spec_change_analysis Configuration/deployment diffs
Table 1: MCP tools for SRE domain context discovery. EoG uses only alert_summary (bootstrap) and get_context_contract (exploration); the ReAct baseline has access to all tools, shell utilities, and can write Python code for data analysis.
For each entity $v$, the Context Packet satisfies the Local Markov Property: it contains entity metadata, 1-hop neighbors, time-windowed observability data, and the current Inbox. Large contexts (>>80K tokens) are chunked with overlap and classified via map-reduce. Domain expertise is encoded in structured prompts organized by phase (bootstrap, explore, finalize), with the core algorithm logic remaining domain-agnostic.
5.2 Experiments
We seek to answer the following research questions:
RQ1.
Does the EoG agent improve majority@$k$ and pass@$k$ over baselines?
RQ2.
Can ReAct agents achieve similar results when the EoG algorithm is explained in the prompt?
RQ3.
How much does SBP improve accuracy beyond the Deterministic Controller alone?
Setup.
We evaluate our methodology on ITBench (Jha et al., 2025), a benchmark of realistic SRE diagnostic tasks on Kubernetes, complementing other agentic task automation benchmarks Shen et al. (2024); Patel et al. (2025); Qiao et al. (2025). We use 35 scenarios spanning misconfigurations, resource exhaustion, and cascading failures. Each scenario is repeated 3 times to measure accuracy and consistency. All experiments use high reasoning effort settings (or equivalent) for each model. We compare EoG against a ReAct baseline; notably, ReAct has access to more tools (all MCP tools plus shell utilities) while EoG uses only two MCP calls per entity, yet controls the investigation structure algorithmically.
RQ1: EoG vs ReAct.
Table 2 compares EoG against ReAct baselines. The results directly address the reliability gap identified in §3.1: while ReAct exhibits large variance between pass@3 and majority@3 (indicating stochastic success), EoG narrows this gap substantially. For example, GPT-5.1 under ReAct achieves only 22.9% pass@3 and 8.6% majority@3 F1 (RC Entity); under EoG, these rise to 88.9% and 86.1% respectively—a near-closed reliability gap.
[…]
We note that the ReAct baseline was provided access to operational topology via tools and the application service architecture in the system prompt—the performance gap is not due to missing structural information. Additionally, ablation experiments removing distributed traces as a data source showed no statistically significant change in results; this is consistent with the redundancy typical of observability data, where the same diagnostic evidence can often be inferred from multiple signals (logs, metrics, events).
Table 2: RQ1: EoG vs ReAct on ITBench (35 scenarios, 3 runs).
  RC Entity RC Reas.
[…]
  ReAct GPT-5.1 22.9 22.9 8.6 11.4 10.0 779K 9.3K
[…]
ReAct GPT-OSS-120B 28.6 28.6 5.7 5.7 8.6 230K 3.8K
[…]
ReAct GPT-OSS-20B
[…]
ReAct Gemini 3 Flash 88.6 94.3 74.3 74.3 72.9 5376K 21.2K
[…]
ReAct Kimi K2 48.6 48.6 28.6 28.6 22.9 1021K 8.8K
[…]
ReAct Minimax M2.1
[…]
ReAct Mistral Large
[…]
  • Best overall shown in bold; other EoG wins in light blue. “–” indicates ReAct baseline could not be run due to non-trivial failure modes with the Codex framework. We omit diagnosis time as it varies due to provider rate limits, reasoning token overhead, and concurrency. Token counts are averaged per trial per scenario and exclude reasoning tokens as providers do not consistently report them.
RQ2: Can ReAct be prompted to follow EoG?
We investigate whether the reliability gains of EoG can be achieved simply by instructing a ReAct agent to follow the EoG algorithm (i.e., “explore neighbors, update beliefs, propagate messages”) in its system prompt, without the external Deterministic Controller. We select a diverse subset of high-performing models due to budget constraints. Table 3 shows the results.
While ReAct+Prompting achieves non-trivial pass@3 recall (40–71%), its majority@3 recall collapses to near zero (0–2.9%). This indicates extreme instability: the agent may occasionally stumble upon the correct diagnosis but cannot reproduce it reliably. The cognitive load of simultaneously managing exploration state, message passing, and causal reasoning overwhelms the model’s instruction-following capacity. This confirms that algorithmic enforcement via a symbolic controller is essential; the EoG protocol cannot be implemented reliably as a prompt-only policy.
Table 3: RQ2: ReAct prompted with EoG algorithm (Recall).
[…]
  GPT-5.1 40.0 2.9
Gemini 3 Flash 71.4 0.0
Kimi K2 42.9 0.0
  • ReAct agent prompted with EoG algorithm instructions. Compare with EoG results in Table 2 (e.g., GPT-5.1: 91.4/88.6 vs. 40.0/2.9 here).
RQ3: SBP Ablation.
Table 4 isolates the contribution of the two architectural components: the Deterministic Controller (graph-guided traversal) and SBP (non-monotonic revision). EoG-SBP uses only the Controller—entities are classified once without revision. Full EoG adds SBP, allowing upstream beliefs to be revised when downstream evidence contradicts initial assessments.
The results show that structured traversal alone (EoG-SBP) captures significant improvement over ReAct, confirming that the exploration pathology (§3.2) is a dominant failure mode. SBP provides substantial additional gains of 7–64% across models, with the largest improvements on weaker models (GPT-OSS-20B: +64.3%) where early misclassifications are more common. This validates the order sensitivity findings from §3.4: while oracle reordering showed 14–70% potential gains from optimal sequencing, SBP recovers a significant portion of this by enabling post-hoc correction without requiring oracle knowledge.
Table 4: RQ3: SBP ablation (maj@3 Recall, RC Entity).
[…]
6 Related Work
Abductive Reasoning with LLMs.
Our work relates to LLM-based investigative agents that perform multi-step reasoning over structured environments Yao et al. (2023); Bei et al. (2025). These systems follow a reason-retrieve-synthesize loop and have been applied to software diagnosis Roy et al. (2024); Xu et al. (2025), cloud operations Chen et al. (2024a); Tian et al. (2025), and graph-augmented QA Sun et al. (2024); Luo et al. (2024). Recent work has also begun to introduce belief-centric abstractions to stabilize LLM agent reasoning by making intermediate assumptions explicit in language Lidayan et al. (2025); Kim et al. (2024); He et al. (2025). These approaches improve consistency by guiding how beliefs are formed or queried, but they treat beliefs as prompt-level artifacts rather than first-class, mutable state. While effective for episodic investigations, they assume monotonic and global updates, and lack explicit mechanisms for belief revision when new evidence contradicts earlier hypotheses.
Table 5: Comparison of LLM/agent baselines.
System Graph as Mutable State Hypotheses Explicit Update Semantics Revision Trigger Propagation Controller Invariants Data Dynamicity Graph Scale
ReAct Yao et al. (2023) append global episodic small
Graph-RAG Edge et al. (2024); Sun et al. (2024) $\circ$ append global summ. episodic medium
Tool-assisted RCA Roy et al. (2024); Xu et al. (2025); Zhang et al. (2025b) $\circ$ $\circ$ append global $\circ$ streaming large
Multi-agent (STRATUS) Chen et al. (2025) $\circ$ $\circ$ append global $\circ$ streaming large
GALA Tian et al. (2025) $\circ$ $\circ$ append global/hybrid $\circ$ streaming large
Ours (EoG) revise msg passing streaming unbounded
Legend: ✓ = yes, ✗ = no, $\circ$ = partial/implicit.
Graph-RAG and Graph-Augmented Agents.
Graph-RAG systems Edge et al. (2024); He et al. (2024); Liu et al. (2025b) extend retrieval to knowledge graphs, enabling multi-hop reasoning and improved coverage over large corpora. Recent work explores optimizing the interface between knowledge graphs and LLMs Marković et al. (2025); Coppolillo (2025) and integrating knowledge graph reasoning into agent workflows Besta et al. (2025). GraphRAG Edge et al. (2024) constructs entity knowledge graphs, partitions them into community hierarchies, and pre-generates summaries aggregated at query time. Think-on-Graph Sun et al. (2024) performs iterative reasoning by exploring knowledge graph relations, while Reasoning on Graphs Luo et al. (2024) generates faithful reasoning plans grounded in graph structure. Theorem-of-Thought Abdaljalil et al. (2025) runs three specialized agents in parallel, stores each agent’s step-wise reasoning as a graph, and then applies Bayesian belief propagation to score these graphs and select the most coherent one. StructGPT Jiang et al. (2023) provides a general framework for reasoning over structured data. However, their inference semantics are append-only: evidence is accumulated globally, with no explicit hypothesis representation, revision trigger, or localized propagation operator. This makes them order-sensitive and brittle under evolving evidence, especially when graphs are large and cannot be fully materialized in context.
Diagnostic Agents in Operations.
LLM agents for diagnosis in cloud operations span several architectures: tool-assisted RCA agents Roy et al. (2024); Xu et al. (2025); Zhang et al. (2025b); Deshpande et al. (2025), multi-agent systems like STRATUS Chen et al. (2025), and graph-augmented workflows like GALA Tian et al. (2025). Recent work also provides systematic taxonomies of failure modes in such diagnostic agents Ma et al. (2025). These systems are explicitly abductive: they generate explanatory hypotheses from partial evidence. Graph-based systems like Groot Wang et al. (2021) and causal approaches like RADICE Tonon et al. (2025) use service dependency graphs to localize faults; similar approaches have been applied to microservice systems Zhu et al. (2024) and IoT environments with dynamic causal graphs Tang et al. (2025). GraphLocator Liu et al. (2025a)-style traversal is captured by our Deterministic Controller’s graph-guided expansion, and GALA/SynergyRCA-style Xiang et al. (2025) evidence aggregation is subsumed by the Context Packet; neither exposes explicit belief revision or localized message passing. Nevertheless, all these systems rely on implicit hypotheses, global summarization, and workflow control rather than controller-enforced invariants over mutable graph state. Even graph-aware workflows remain monotonic, without first-class revision semantics or message passing that localizes updates to affected neighborhoods. Our ReAct baseline subsumes the inference semantics of these approaches: it has access to all diagnostic tools (topology queries, log/metric analysis, trace inspection) plus shell utilities and code execution, strictly more capability than published RCA agents. Yet even the best-performing ReAct configuration (Gemini 3 Flash) achieves only 74% majority@3 F1 on ITBench, compared to 92% for EoG with the same model. This 18-point gap persists across model families, confirming that gains arise from state semantics (explicit beliefs, revision, propagation), not domain engineering or tool access.
Causal Inference.
Causal discovery and effect estimation Wan et al. (2024); Vowels et al. (2022); Yao et al. (2021) target different outputs (DAGs, treatment effects) under static, fully-observed data; they do not model revisable explanatory hypotheses under streaming partial evidence and are thus not comparable to abductive graph reasoning for baseline consideration.
Putting EoG in perspective.
Our contribution is an inference abstraction for abductive reasoning under streaming evidence and unbounded graphs, where global context is infeasible and explanations must be revised online. The baselines above cover the strongest LLM instantiations of investigative reasoning; differences arise from state semantics, not domain or prompting. Table 5 summarizes the comparison across inference-semantic axes. We compare against representative ReAct, Graph-RAG, and diagnostic agents, and ablate revision and propagation—demonstrating that gains arise from state semantics rather than domain-specific engineering.
7 Conclusion
We framed operational diagnosis as abductive inference over graphs and introduced Semantic Belief Propagation, which separates local LLM reasoning from deterministic control to enable non-monotonic belief revision. Experiments on ITBench show gains in accuracy, stability, and token efficiency over ReAct-style agents. This architecture yields reproducible, auditable explanations under realistic production constraints.
Appendix A ReAct Agent Implementation Details
We provide implementation details for the ReAct baseline agent Yao et al. (2023) used in our evaluation. The baseline leverages OpenAI Codex CLI (v0.76) OpenAI (2025), an open-source coding agent that runs locally and supports extensible tool integration via the Model Context Protocol (MCP).
A.1 OpenAI Codex CLI Architecture
Codex CLI CoreOpenAI Codex CLI (v0.76)File OpsShellPythonSummarizeModel Context Protocol (MCP)SRE ToolsCustom MCPLLM LoopBuilt-inInterfaceExtensions
Figure 7: OpenAI Codex CLI architecture. The core agent loop interfaces with built-in tools through the Model Context Protocol. Domain-specific MCP servers extend capabilities.
Codex CLI implements a disaggregated architecture where the LLM reasoning loop is separated from tool execution OpenAI (2025). The core is written in Rust (codex-rs) and provides:
  • Sandboxed Execution: Shell and code execution occur in isolated subprocesses with configurable permission levels (read-only, write workspace, full access). The sandbox restricts filesystem writes, network access, and system modifications by default.
  • MCP Tool Servers: Tools run as separate processes communicating via the Model Context Protocol. This separation enables: (1) crash isolation—a failing tool does not crash the agent; (2) resource limits—each tool server can be independently constrained; (3) extensibility—domain tools are added without modifying the core.
  • Automatic Context Summarization: When tool outputs exceed context limits, Codex summarizes content to preserve important information while staying within token bounds.
  • State Persistence: Python execution maintains state across invocations; file operations are transactional within the workspace.
A.2 ReAct Agent Prompt
The ReAct baseline operates as a Site Reliability Engineering (SRE) support agent designed for offline incident analysis. The agent follows a structured investigation workflow with explicit phases for context discovery, symptom analysis, hypothesis generation, evidence collection, and causal chain construction. The full system prompt is provided below, with environment variables (\$SNAPSHOT_DIRS, \$WORKSPACE_DIR) populated at runtime.
ReAct with Code System Prompt Task: You are an expert SRE (Site Reliability Engineer) and Kubernetes SRE Support Agent investigating a production incident from OFFLINE snapshot data. You are a highly capable tool-using agent able to: Diagnose Kubernetes failures; Correlate alerts, events, traces, and metrics; Identify contributing factors (not just root cause); Perform data analysis using Python when useful. INCIDENT SNAPSHOT DATA LOCATION
Your incident data is located in \$SNAPSHOT_DIRS (READ-ONLY). Do NOT search the filesystem for additional data. Start by listing the contents of these directories. Working directory: \$WORKSPACE_DIR. Application topology available at \$WORKSPACE_DIR/app.json.
FINAL OUTPUT FORMAT (MANDATORY)
Generate a JSON diagnosis identifying all Kubernetes entities: entities that CAUSED the incident (contributing_factor=true), entities IMPACTED but did not cause it (contributing_factor=false), and the propagation chain showing how the incident spread. Requirements: explain all firing alerts, provide reasoning/evidence for every entity, construct fault propagation chain, use Python for data analysis when necessary. Write diagnosis to \$WORKSPACE_DIR/agent_output.json. Validate JSON using jq after writing.
ENTITY NAMING CONVENTION
All entities MUST use format: namespace/Kind/name. Examples: otel-demo/Deployment/ad, otel-demo/Service/frontend. Do NOT include UIDs.
PROPAGATION CHAIN (MANDATORY)
Construct chain: Root Cause $\to$ Intermediate Services $\to$ Impacted Services. Each link includes: source (causing entity), target (affected entity), condition (what state caused propagation), effect (observable effect on target).
RULES FOR CONTRIBUTING FACTORS
contributing_factor=true (IRREDUCIBLE): Mark ONLY if entity is an independent cause not fully explained by another contributing factor. Use “irreducibility test”: if failure is explained by “upstream X failed” with propagation edge $X\to$ entity, then entity is NOT irreducible $\to$ set false.
contributing_factor=false (DERIVED): Mark entities that are downstream effects/symptoms caused by another contributing factor. IMPORTANT: Do NOT mark both cause and derived symptom as true. Multiple contributing_factors allowed ONLY if truly independent. DEBUGGING PRINCIPLES
(1) Differential Observability: Compare replicas and time windows. (2) Occam’s Razor: Choose simplest explanation consistent with evidence. (3) Duration Matching: Theory must explain entire incident duration. (4) Follow Breadcrumbs: Let alerts/errors guide investigation. (5) No Jumping to Conclusions: Validate every hypothesis with evidence. (6) Chaos Files $\neq$ Active Chaos: Verify experiment was running AND time-aligned. (7) Semantic Name Normalization: Try variations, strip suffixes, search partial matches.
INVESTIGATION WORKFLOW (DO NOT SKIP)
Phase 1 – Context Discovery:
List available files; read topology.md to understand dependencies.
Phase 2 – Symptom Analysis: Read all alert files; compute start/end time, duration, frequency; build Alerts Table. Phase 3 – Hypothesis Generation: Create initial hypotheses; create validation plan for each. Phase 4 – Evidence Collection Loop: Use tools and Python to gather log/event/metrics/trace evidence; validate/refute hypotheses; explain firing alerts. Phase 5 – Causal Chain Construction: Build chain: [Config Error] $\to$ [CrashLoop] $\to$ [Service Down] $\to$ [Frontend 5xx]. Phase 6 – Conclusion: Ensure all alerts explained, all entities included, contributions correctly labeled; output final JSON. PYTHON ANALYSIS (ENCOURAGED)
Write Python snippets to: parse alerts/events/metrics/traces, join datasets by timestamp or UID, extract failing spans, compute metrics deltas, identify patterns (CrashLoopBackOff, OOMKilled, 5xx spikes). Persist snippets in Python files.
PROHIBITED ACTIONS
NEVER read ground_truth.yaml; NEVER output anything other than final JSON; NEVER hallucinate K8s objects; NEVER leave alerts unexplained.
A.3 Expected Output Schema
Both the ReAct baseline and EoG agent must produce a structured JSON diagnosis with three key components:
"entities": [{
"name": "namespace/Kind/name",
"contributing_factor": true|false,
"reasoning": "Explanation for involvement",
"evidence": "Supporting facts summary"
"propagations": [{
"source": "namespace/Kind/source-name",
"target": "namespace/Kind/target-name",
"condition": "What caused propagation",
"effect": "What effect was observed"
"alerts_explained": [{
"alert": "<alert name>",
"explanation": "Why the alert fired",
"explained": true|false
A.4 SRE Domain Tools (MCP Server)
We implement a custom MCP server providing specialized SRE investigation tools. Table 6 provides detailed function signatures for the primary tools.
Table 6: SRE Utility Tools Available via MCP
Tool Description
build_topology Build operational topology graph from application architecture and Kubernetes objects
topology_analysis Analyze topology for dependencies, service context, infrastructure hierarchy
metric_analysis Analyze metrics with filtering, grouping, derived metrics (SQL-like interface)
get_metric_anomalies Detect anomalies in time-series metrics
event_analysis Analyze K8s events with filtering, grouping, aggregation
get_trace_error_tree Analyze distributed traces with error trees and latency percentiles
alert_summary High-level alert overview: type, entity, duration, frequency
alert_analysis Detailed alert analysis with filtering and grouping
k8s_spec_change_analysis Track K8s object spec changes over time
get_k8_spec Retrieve K8s spec for a specific resource
get_context_contract Aggregated context for an entity (events, alerts, traces, metrics, spec changes)
Key Tool Signatures.
metric_analysis: SQL-like interface for metric analysis supporting batch queries, derived metrics, and flexible grouping.
def metric_analysis(
base_dir: str, # Metrics directory path
k8_object_name: str = None, # Optional: specific object
object_pattern: str = "*", # Glob pattern (e.g., ’pod/*’)
metric_names: List[str] = [], # Metrics to analyze
eval: str = None, # Derived metric expression
group_by: str = None, # Grouping column
agg: str = "mean", # Aggregation function
start_time: str = None, # Time range start
end_time: str = None # Time range end
) -> dict
get_trace_error_tree: Analyzes distributed traces to find critical paths with regressions.
def get_trace_error_tree(
trace_file: str, # Path to traces TSV
service_name: str = None, # Filter by service
pivot_time: str = None, # Timestamp for before/after
delta_time: str = "5m", # Comparison window size
error_threshold_pct: float = 10, # Error change threshold
latency_threshold_pct: float = 10 # Latency change threshold
) -> dict
get_context_contract: The primary tool for EoG, returning full operational context for an entity.
def get_context_contract(
k8_object: str, # K8s object (Kind/name format)
snapshot_dir: str, # Snapshot directory path
topology_file: str = None, # Optional topology JSON path
start_time: str = None, # Time range start
end_time: str = None, # Time range end
page: int = 1, # Page number
deps_per_page: int = 3 # Dependencies per page
) -> dict
Appendix B EoG Implementation Details
This section provides complete architectural and prompt specifications for the EoG (Explanations over Graphs) agent, sufficient for reproduction.
B.1 Architecture Overview
Figure 8 illustrates the complete EoG execution flow, from initialization through the event-driven inference loop to final output generation.
[…]
Figure 8: EoG implementation architecture. Initialization: Multi-phase context acquisition filters events and spec changes, then bootstraps initial candidates. Event Loop: Each Activate event triggers context fetch, policy inference, and conditional belief broadcast. The loop continues until the queue empties or budget exhausts. Output: The frontier $\mathcal{F}$ is computed and a structured report is generated.
B.2 Event-Driven Architecture
The implementation realizes the ActiveSet abstraction from §4.3 as an event-driven message queue. This architectural choice enables concurrent node evaluation without changing the algorithm’s semantics—each event type corresponds directly to operations in the abstract algorithm.
Table 7: Event Types and Semantics
Event Semantics
Activate(v, reason) Evaluate node $v$; equivalent to ActiveSet.pop(v)
Message(u, v, $B_{u}$, t) Deliver belief $B_{u}$ to $v$’s inbox
BeliefChanged(v) Broadcast $B_{v}$ to neighbors; triggers re-activation
The event queue processes Activate events by: (1) checking visit eligibility, (2) collecting inbox messages, (3) fetching local context, (4) invoking the abductive policy, (5) applying damping if oscillating, (6) emitting BeliefChanged if belief changed, and (7) emitting Activate for discovered or proposed nodes. This is semantically equivalent to the ActiveSet-based loop but enables the actor model for concurrent processing.
B.3 Termination Safeguards
Beyond the damping heuristic ($k_{\text{thresh}}$) described in §4.3, the implementation adds visit limits for robustness:
  • Max Visits: Each node may be evaluated at most $k_{\text{max}}$ times (default: 5), regardless of belief stability.
  • Cooldown: Re-evaluation requires at least $k_{\text{cool}}$ other evaluations since last visit (default: 2), preventing rapid re-activation loops.
Together with damping, these bound total inference to $O(\min(k_{\text{thresh}},k_{\text{max}})\cdot|V_{S}|)$.
B.4 Checkpointing
The implementation supports fault-tolerant checkpointing at quiescent points between event processing. Checkpoints capture node beliefs, inbox contents, pending events, and the explanatory graph $(V_{S},E_{S})$, enabling recovery without restarting the investigation.
B.5 Initialization Prompts
The investigation begins with a multi-phase initialization that acquires bounded context and selects initial candidates.
B.5.1 Investigation Window Selection
The first step determines the temporal scope by identifying “golden signal” alerts.
select_investigation_window You are an SRE agent. Your goal is to choose an investigation time window [start_time, end_time] that best captures the onset and peak of the incident, using golden signal alerts as anchors. Golden signals include alerts indicating: Errors (request error rate, 5xx), Latency (p95/p99), Traffic (throughput drop, no requests). Ignore noisy platform alerts (watchdog, inhibitors) unless they directly explain golden-signal symptoms. Input: Alert Summary (JSON with names, entities, timestamps, severity), Application Architecture (service dependencies). Output: JSON with start_time, end_time, anchor_alerts, golden_signal_alerts, reasoning. Rules: Prefer application services over infrastructure alerts. start_time $\leq$ end_time. Use timestamps from alert summary.
B.5.2 Event Filtering
Kubernetes clusters generate thousands of events per hour. We filter events in pages, selecting those relevant to the investigation.
filter_events You are filtering Kubernetes events to identify those relevant to an ongoing incident investigation. Relevance Criteria --- INCLUDE: Warning/Error type events Events on alerting entities or their dependencies Pod lifecycle failures (CrashLoopBackOff, OOMKilled, Evicted) Scheduling failures (FailedScheduling, FailedMount) Probe failures (Unhealthy, ProbeWarning) Resource issues (FailedCreate, QuotaExceeded) Deployments/restarts preceding alerts EXCLUDE: Normal successful operations (Scheduled, Started, Created) unless systematic churn or temporally preceding alerts. Prioritization: Return at most {{max_per_page}} events. Rank by: (1) events preceding alerts, (2) events on alerting entities, (3) failure events, (4) config changes. Prefer semantic diversity---avoid redundant events conveying the same information. Output: JSON with relevant_events array, total_relevant_before_limit, reasoning.
B.5.3 Spec Change Filtering
Configuration changes are among the most common root causes. We filter spec changes with similar pagination.
filter_spec_changes You are filtering Kubernetes spec changes to identify those relevant to an ongoing incident investigation. Configuration changes are one of the biggest failure causes. Prioritization Guidance: 1. Highest: Changes just before alerts (within minutes)---prime root cause suspects 2. Medium: Changes during incident window 3. Lower: Changes well before/after incident Key Question: ‘‘Did this change happen before the symptoms started?’’ If yes, strong candidate for root cause. Output: JSON with relevant_changes array (max {{max_per_page}}), total_relevant_before_limit, reasoning.
B.5.4 Candidate Selection (Bootstrap)
With filtered events and spec changes, the bootstrap prompt selects initial investigation candidates.
bootstrap You are an SRE agent. Review active alerts and select initial candidate entities for investigation. Strategy (Golden Signal Priority): 1. Identify minimum set of alerts that cannot be explained by other alerts/events 2. Ignore positive goodput/watchdog alerts 3. Infrastructure alerts (OOM, throttling) help explain golden signals Strict Ordering by Causal Evidence Strength: STRONGEST (FIRST): Entity with spec change preceding alert start STRONG: K8s event (OOMKilled, CrashLoopBackOff) on alerting entity MEDIUM: Change on entity with known dependency relationship WEAKEST: Alerting service with NO associated change---likely symptom Causal Path Requirement: A spec change is only relevant if there’s a causal path to the alerting application (dependency, traffic source, config, fault injection). Output: JSON with initial_candidates array (node_id, start_time, end_time, alert_type, reason_for_selection, identified_facts, identified_events), reasoning.
B.6 Abductive Policy Prompt
The explore prompt implements local inference, reasoning over both local evidence and neighbor messages.
explore (Abductive Policy $\pi_{\text{abd}}$) You are an SRE support engineer doing RCA using the EoG algorithm. Objective: Help the controller converge to a small root-cause frontier. Do NOT enumerate all impacted entities. Definitions: Healthy (L=0): Operating normally in the alert window Symptom (L=0): Degraded, but evidence points to upstream cause Origin (L=1): Has citable change (config, deployment, resource) that temporally precedes and explains incident Defer (L=2): Evidence inconclusive Messages from Neighbors (Inbox): When neighbor claims SYMPTOM, consider if YOUR component caused their failure. When neighbor claims ORIGIN, consider if YOUR issues are symptoms of their fault. Core Investigation Principles: 1. ‘‘What Changed’’ Discipline: A node cannot be Origin if nothing changed within it 2. Transitive Traffic Amplification: Resource exhaustion without local change $\to$ suspect traffic surge 3. Differential Observability: Compare replicas, time windows 4. Occam’s Razor + Multiple Causes: Prefer unified explanation, accept independent causes when evidence supports Hard Rules: Evidence required for every decision (cite logs/events/metrics/traces) Never revisit entities in Visited set Max 2 next_candidates per step One direction at a time: upstream $|$ downstream $|$ ownership $|$ infrastructure Output: JSON with direction, local_origin (Healthy||Symptom||Origin||Defer), reasoning, propagations array, next_candidates array.
B.7 Finalize Prompt
When the event queue empties or budget exhausts, the finalize prompt synthesizes findings.
finalize Summarize the investigation, identifying Root Causes (Frontier) and Causal Graph ($G_{S}$). If no frontier found, make the most educated guess given all evidence. Critical Distinction: Root Cause (contributing_factor: true) = Entity that CAUSED the problem: config changes, deployments, infrastructure failures, fault injections, traffic pattern changes Symptom (contributing_factor: false) = Entity that EXHIBITED the problem: elevated error rates, failures due to upstream causes Validation Checkpoint: For each contributing_factor: true entity, answer: (1) What specifically changed? (2) When did it change? (3) How does this change explain the symptoms? If you cannot answer all three, it’s likely a symptom. Contributing Factor Rules: True (Irreducible): Origin in Frontier with no incoming causal edge from another Origin False (Derived): Symptom affected by propagation; failure explained by upstream Output: JSON with entities array (name, contributing_factor, reasoning, evidence), propagations array (source, target, condition, effect), alerts_explained array.
B.8 Implementation Notes
The system comprises approximately 3,500 lines of Rust with prompts maintained as Handlebars templates. Key implementation choices:
  • Temperature Zero: All LLM calls use temperature 0 for determinism.
  • Structured Output: All prompts request JSON validated against schemas. Invalid responses trigger a single retry with error correction.
  • Token Budgeting: The Context Contract implements token-aware pagination. When filtered events/changes exceed budgets, tournament-style reduction selects the most relevant items.
  • Entity Normalization: All prompts mandate namespace/Kind/name format for consistent deduplication and graph construction.
The implementation is available as part of the open-source Codex project OpenAI (2025).
Appendix C LLM-as-a-Judge Evaluation Framework
We employ an LLM-as-a-Judge (LAAJ) evaluation framework to assess agent outputs against ground truth root cause analysis. The judge evaluates two primary metrics—Root Cause Entity Identification (Precision, Recall, F1) and Root Cause Reasoning Accuracy—in a single LLM call.
C.1 Entity Mapping Challenge
To understand why entity normalization and mapping is critical, consider Scenario-1 from our benchmark involving a traffic overload incident.
C.1.1 Ground Truth Structure
The ground truth defines entities using semantic identifiers with regex-based filters:
Listing 1: Ground Truth Entity Groups (Scenario-1)
groups:
- id: load-generator-pod-1
kind: Pod
filter: [load-generator-.*]
namespace: otel-demo
root_cause: true
- id: load-generator-deployment-1
kind: Deployment
[…]
namespace: otel-demo
root_cause: true
- id: flagd-config-1
kind: ConfigMap
[…]
namespace: otel-demo
root_cause: true
- id: frontend-proxy-service-1
kind: Service
[…]
namespace: otel-demo
C.1.2 Alias Groups
Multiple Kubernetes objects represent the same logical service. The ground truth captures this through alias groups:
Listing 2: Ground Truth Aliases
aliases:
- [load-generator-service-1, load-generator-pod-1,
load-generator-deployment-1]
- [frontend-proxy-service-1, frontend-proxy-pod-1]
If a model identifies otel-demo/Service/load-generator as a root cause, it receives credit because it is an alias of load-generator-pod-1.
C.1.3 Propagation Chain
The ground truth defines the causal chain showing how the fault propagated:
Listing 3: Ground Truth Propagation Chain
propagations:
- source: flagd-config-1
target: load-generator-deployment-1
condition: flagd-config updated to enable higher traffic
effect: load-generator deployment applies new config
- source: load-generator-deployment-1
target: load-generator-pod-1
condition: updated deployment rolls out pods
effect: load-generator pod configured with higher users
- source: load-generator-pod-1
target: load-generator-service-1
condition: pod generates high number of requests
effect: service receives high request volume
- source: load-generator-service-1
target: frontend-proxy-service-1
condition: frontend-proxy overloaded
effect: Error rate and latency above threshold
C.1.4 The Mapping Challenge
Actual Kubernetes object names include generated suffixes. The evaluation must map agent-predicted entities to ground truth semantic IDs using regex filters. For example:
  • load-generator-pod-1 matches otel-demo/Pod/load-generator-c7cbc4c99-xyz12
  • flagd-config-1 matches otel-demo/ConfigMap/flagd-config
  • load-generator-deployment-1 matches otel-demo/Deployment/load-generator
C.2 Judge System Prompt
The LAAJ system prompt defines the evaluation protocol:
LAAJ System Prompt You are an expert AI evaluator specializing in Root Cause Analysis (RCA) for complex software systems. Your role is to act as a fair, consistent, and precise judge when evaluating RCA model outputs. You will be provided with: 1. A Ground Truth (GT) JSON object: true root cause(s), failure propagation path(s), and entity definitions. 2. A Generated Response JSON object: output from an RCA model to evaluate. Phase 1: Normalization and Entity Mapping
Before any scoring, normalize and map entities from Generated Response to Ground Truth based on explicit evidence from entity metadata. Do not infer mappings based on position in propagation chain.
Definition of a Confident Match:
A Generated Response entity is a confident match ONLY IF its name field clearly corresponds to the filter and kind of the GT entity.
Entity Name Format:
Generated Response entities use format: namespace/Kind/name
Examples: otel-demo/Deployment/flagd, otel-demo/Service/frontend
CRITICAL - Alias Handling:
If a predicted entity matches an ALIAS of a root cause entity, it MUST be counted as correct. The GT.aliases field contains arrays of equivalent entity IDs. If ANY entity in an alias group has root_cause: true, then matching ANY entity in that alias group counts as correctly identifying the root cause.
[…]
C.3 Evaluation Metrics
The LAAJ framework evaluates two primary metrics:
  • Root Cause Entity (Precision, Recall, F1): Match contributing_factor=true entities against GT root causes with alias support.
  • Root Cause Reasoning (Average): Semantic similarity of condition descriptions using 0/0.5/1 scoring.
C.3.1 Root Cause Entity Identification
Root Cause Entity Scoring Method: 1. Extract all entities with root_cause: true in GT.groups. 2. Handle Aliases: Check GT.aliases for any alias groups containing a root cause entity. ALL entities in such groups are valid root cause matches. 3. Extract all entities with contributing_factor: true from Generated Response, preserving order. 4. Normalize entities to compare against GT root causes. 5. For EACH predicted entity (in order), determine if it matches any GT root cause entity OR any alias. 6. Output per-entity match list showing which predicted entities matched GT entities. 7. Calculate Recall = Correct / Total GT (count each alias group only once). 8. Calculate Precision = Correct / Total Predicted. 9. Calculate F1 = (2 $\times$ Recall $\times$ Precision) / (Recall + Precision). Output: List of GT entities, ordered list of predicted entities with per-entity match boolean, Precision, Recall, and F1.
C.3.2 Root Cause Reasoning Accuracy
Root Cause Reasoning Scoring Method: 1. Extract the condition of every correct root cause entity from Generated Response. 2. Extract the condition of corresponding entities from GT. 3. For each correct root cause entity, compare conditions: Score 1.0: Model’s reason is semantically equivalent to GT’s condition Score 0.5: Model’s condition describes similar resource/action but at least one key detail is missing or imprecise (e.g., GT=‘‘configmap Y featureflag X set’’, Model=‘‘configmap updated’’) Score 0.0: Incorrect or absent model condition 4. Calculate average of all entity reasoning scores. 5. If no root cause entity correctly identified, score = 0. Output: Average reasoning score.
C.4 Evaluation Prompt Template
Evaluation User Prompt Given the following Ground Truth (GT) and Generated Response, evaluate the response according to the scoring rubric. ## Ground Truth (GT):
{ground_truth}
## Generated Response:
{generated_response}
## Task: 1. Perform Phase 1: Normalize and map entities. 2. Calculate ROOT_CAUSE_ENTITY metrics (Precision, Recall, F1). 3. Calculate ROOT_CAUSE_REASONING score. 4. IMPORTANT: Use calculator_tool for ALL mathematical calculations. CRITICAL: Final response MUST be ONLY the JSON object with scores.